Question 4 2 pts Find P(-1.47< z< 1.79) = places. Round to 4 decimal
Answer #1

We have to find P(-1.47 < z < 1.79). We know that P(-1.47 < z < 1.79) can be written as P(z < 1.79) - P(z < -1.47).Thus,

P(-1.47 < z < 1.79) = P(z < 1.79) - P(z < -1.47) = 0.9633 - 0.0708 = 0.8925(rounded up to four decimal places). [From normal distribution table]

Thus, P(-1.47 < z < 1.79) = 0.8925 .